The Lowly Engineer
  • Metasys

    Gen Test, Challenges the OSC

    copy of email to Fred:

     

    Fred,

    Many problems at OSC today.  You guessed it Gen Test!  When they transferred back they took out a breaker for the 120V transformer that serves OSC.  Everything went down.  Now as the Metasys devices come back up, we are discovering everything moving way to slow.  All the OR reheat valves move very slow and one of the biggest problems of the morning was the steam valve for the heat exchanger.  It would not move but a percent every five minutes, so we lost all the OR’s and they had to delay Surgeries.   We got it all back up with overrides and monitoring.

     

    Do you think Kyle’s changes got blown out?   I also noticed that there was no reset schedule in, it was at 120* high and low.  I put the “output Low at 140” to try and help the temps come up with a little warmer water. (I actually had it at 160* but we overshot a couple rooms, so I bumped back down).

  • Metasys

    OSC ATU 53-4 Service Core C RM 1954 Steam visible at supply registers

    The steam supply serving ATU 53-4 has been valved off  in M.E.R. 32.  OSC staff are seeing steam at the supply registers under certain conditions.  JCI will adjust programming next Tuesday morning prior to cases starting.  Access to the supply (ball valve) shut off requires a tall ladder, currently placed underneath valve…  Will restore once new limit set points are established.

  • Metasys

    North Tower RM 259 VMA 2-27

    Per Dr. Vely the patient in North Tower RM 259 (burn victim) is to be kept exceptionally warm.  Zone temperature currently at 78.7 through overrides is okay.  Annotations placed on the overrides as a local reminder…  Release or manipulate as needed.

  • Metasys

    AC27-10 & AC28-10

    The final programs have been downloaded & tested. The humidfiers have control panel issues & will be dealt with next week. I’ve mapped 27 & 28 into the user view.